OpenStudy (anonymous):

Given a cylinder with a surface area of 60 pi units squared and a height of 7 units. Find the volume of a sphere that has the same size radius as the given cylinder. Show all work.

OpenStudy (anonymous):

What is the formula for the surface area of a cylinder?

OpenStudy (anonymous):

Surface Area of a Cylinder = 2 pi r squared + 2 pi r h

OpenStudy (anonymous):

Ok, and what is the formula for the volume of a sphere?

OpenStudy (anonymous):

V=4/3pi(r^3)

OpenStudy (anonymous):

So now, solve the surface area formula for r (in terms of the area). Then plug in the area they gave you to find r. Then plug that r into the volume equation to find V.

OpenStudy (anonymous):

Surface Area of a Cylinder = 2 pi r squared + 2 pi r h 60 pi =2 pi r ^2+2 pi r *7 divide whole equation with 2pi and rearrange r^2+7r-30=0 (r+10)(r-3)=0 r=3 as r can't be negetive substitute r to find out volume of sphere in V=4/3pi(r^3) V=4/3*3.14*27=113.04=36pi cubic units

OpenStudy (anonymous):

@_@ I Just Finished The Question When I Got The Reply o.o (Lol Fail) And I Got 113.09 units cubed.. Close Enough (: THANKS YOU<3